If the effective mass of electron in copper is m* = 1.01 me (me =



mass of an electron) and the electric conductivity of copper is 5.76 x107-
1
Expert's answer
2021-09-24T09:26:45-0400

Find the electron concentration number density:


"n=\frac{N_A}{\Mu}\rho=8.43·10^{28}\text{ m}^{-3}."

The relaxation time is


"\tau=\frac{m^*\sigma}{e^2n}=2.45·10^{-14}\text{ s}."